# Enhanced Partially Relevant Video Retrieval through Inter- and Intra-Sample Analysis with Coherence Prediction

### v2

@@ 4.1.2. Evaluation Metrics
Following previous works (Dong et al., 2022a, 2023; Wang et al., 2024b), we adopt the rank-based metrics Recall Rate at $k$ ($R@k$) and the sum of all recall rates ($SumR$). $R@k$ measures the proportion of relevant items successfully retrieved within the top-$k$ results, with $k$ set to {1, 5, 10, 100}. Higher recall rates indicate better retrieval accuracy.
4.1.3. Implementation Details
@@
For fair comparison with prior works on PRVR (Dong et al., 2022a; Jiang et al., 2023; Wang et al., 2024b; Yin et al., 2024; Dong et al., 2023; Zhang et al., 2025; Moon et al., 2025; Cho et al., 2025; Jun et al., 2025), we utilize the same visual and textual features provided by (Dong et al., 2022a). The visual and textual features are extracted by pre-trained I3D (Carreira and Zisserman, 2017) with ResNet152 (He et al., 2016) and RoBERTa (Liu et al., 2019). The feature dimension $D$ is set to 384. $\lambda_{1}$, $\lambda_{2}$, $T_{m}$, $m$, and $\alpha$ are set to 0.02, 0.04, 32, 0.2, and 0.7 following (Dong et al., 2022a; Wang et al., 2024b). $g$ is set to 8. For model training, we utilize the Adam optimizer (Kingma and Ba, 2015) with an initial learning rate of 2.5e-4 and a mini-batch size of 128. The model is trained for 100 epochs with an early stopping strategy. All experiments are conducted on a single NVIDIA RTX 3090 GPU. More implementation details are provided in the appendix.
4.2. Comparison with State-of-the-Arts
4.2.1. Compared Methods
We compare the proposed method with other state-of-the-art methods on TVR, ActivityNet Captions, and Charades-STA. These methods include PRVR (Dong et al., 2022a; Jiang et al., 2023; Wang et al., 2024b; Yin et al., 2024; Dong et al., 2023; Zhang et al., 2025; Moon et al., 2025; Cho et al., 2025; Jun et al., 2025), T2VR (Dong et al., 2021, 2022b; Wu et al., 2023; Li et al., 2023), and large multimodal models (Wang et al., 2024a). The results are sourced from (Dong et al., 2022a; Wang et al., 2024b) and corresponding papers.
4.2.2. Performance
@@
As summarized in Table 1, our method achieves state-of-the-art (SOTA) results on all three datasets across all metrics. Notably, we outperform the previous SOTA method MGAKD (Zhang et al., 2025) without using additional CLIP models. Besides, we also surpass the Large Multimodal Model (LMM) InternVideo2 (Wang et al., 2024a) by a significant margin. InternVideo2 merely considers the similarity between text queries and entire videos, neglecting the essential fine-grained understanding of moment events. In contrast, our method fully leverages the rich semantics of distinct moment events and learns to distinguish them. Comparison highlights the advantages and expertise of our method in the PRVR task.
4.2.3. Efficiency
@@
We evaluate the efficiency of our proposed model and existing models in the inference phase. For fair comparison, all experiments are conducted using the same batch size and a single RTX 3090 GPU on ActivityNet Captions. As shown in Table 2, with the same backbone and fewer learnable parameters, we achieve faster inference speed and better retrieval accuracy. These results demonstrate the effectiveness and efficiency of our proposed model.
@@
Table 2. Efficiency comparison on ActivityNet Captions. All experiments are conducted under identical conditions.
Method Backbone Model Size Run-Time SumR
MS-SL (Dong et al., 2022a) I3D + RoBERTa 3.37M 146.63s 140.1
GMMFormer (Wang et al., 2024b) I3D + RoBERTa 11.37M 87.69s 146.0
Ours I3D + RoBERTa 4.65M 83.01s 160.4
4.3. Ablation Studies
4.3.1. Main Ablation Studies
Table 3. Main ablation studies of the proposed key modules on Charades-STA. The best results are in bold.
Row Setting R@1 R@5 R@10 R@100 SumR
ICE IRM TCP
1 1.5 6.8 11.3 49.5 69.1
2 2.2 7.5 12.4 52.2 74.3
3 2.1 7.8 12.9 51.9 74.7
4 1.7 7.0 12.0 50.5 71.2
5 2.4 8.3 13.8 52.1 76.6
6 2.5 8.5 13.6 52.2 76.8
7 2.6 8.8 14.2 52.6 78.2
8 2.9 9.2 14.9 54.3 81.3
Table 4. Ablation study of loss terms in the Intra Redundancy Mining module on Charades-STA.
Row Loss Terms R@1 R@5 R@10 R@100 SumR
$\mathcal{L}_{neg}$ $\mathcal{L}_{red}$
1 2.5 8.5 13.6 52.2 76.8
2 2.3 8.9 14.9 53.2 79.3
3 2.3 8.7 14.7 52.8 78.5
4 2.9 9.2 14.9 54.3 81.3
@@ 4.3.1. Main Ablation Studies
(1) Each module independently improves performance over the base model (Row 1-4), suggesting their effectiveness on refining the cross-modal semantic space.
@@
(3) Integrating all three modules leads to the best performance (Row 8), with significant gains across all metrics. In particular, TCP further improves ICE and IRM. ICE and IRM alone achieve a 9.1 absolute improvement in SumR (Row 1 and 7, 69.1 → 78.2). When combined with TCP, the performance gain increases to 10.1 (Row 4 and 8, 71.2 → 81.3).
4.3.2. Intra Redundancy Mining
@@
We verify the effectiveness of each loss function in the Intra Redundancy Mining (IRM) module. As shown in Table 4, both $\mathcal{L}_{neg}$ (Row 2) and $\mathcal{L}_{red}$ (Row 3) lead to boosts on SumR. Moreover, combining $\mathcal{L}_{neg}$ and $\mathcal{L}_{red}$ (Row 4) yields the highest SumR. The effectiveness arises because $\mathcal{L}_{neg}$ directly suppresses redundancy by pushing away redundant and query-relevant moments, while $\mathcal{L}_{red}$ implicitly strengthens cross-modal alignment by correlating video-view and query-view redundancies.
Table 5. Ablations of the group number $g$ in the metric SumR.
Row $g$ Charades-STA TVR ActivityNet Captions
1 4 80.0 187.5 156.1
2 8 81.3 194.0 160.4
3 16 78.2 184.0 152.1
4 32 75.1 180.1 148.6
Table 6. Ablation study of the Temporal Coherence Prediction module on Charades-STA.
Row Setting R@1 R@5 R@10 R@100 SumR
Video Moment
1 2.6 8.8 14.2 52.6 78.2
2 2.7 9.0 14.6 53.4 79.7
3 2.8 8.9 14.8 53.7 80.2
4 2.9 9.2 14.9 54.3 81.3
Table 7. Effect on different base models on Charades-STA. We directly apply our proposed three modules to existing PRVR methods.
Method Setting R@1 R@5 R@10 R@100 SumR
MS-SL Origin 1.8 7.1 11.8 47.7 68.4
+ Ours 2.4 9.1 14.6 53.9 80.0
GMMFormer Origin 2.1 7.8 12.5 50.6 72.9
+ Ours 3.0 9.5 15.3 55.1 82.9
@@
Figure 5. Visualization comparisons of retrieval results between our method, MS-SL (Dong et al., 2022a), and GMMFormer (Wang et al., 2024b) on ActivityNet Captions. Top-5 retrieval results are shown from left to right. Ground-truth videos are marked with red boxes.
(a) Base Model
(b) Full Model (Ours)
@@
Figure 6. t-SNE visualization (van der Maaten and Hinton, 2008) of text and video features on Charades-STA. (a) is the base model trained with only $\mathcal{L}_{base}$. (b) shows the full model trained with the full setup.
@@
Figure 7. Text-moment similarity on TVR. The base model is trained with only $\mathcal{L}_{base}$, and the full model is trained with the full setup. We smooth out similarity intervals for better readability. Zoom in for better visibility.
4.3.3. Temporal Coherence Prediction
@@
We assess the group number $g$ in the Temporal Coherence Prediction (TCP) module, which controls the granularity of temporal coherence learning. As shown in Table 5, performances decline as $g$ ¿ 8 in all datasets due to increased task difficulty. Notably, HERO’s method (Li et al., 2020) that assigns unique labels (i.e., $g$ = 32 for $T_{m}$ = 32 moment sequences) obtains the lowest results. The optimal results are achieved at $g$ = 8, suggesting that moderate grouping balances task difficulty and learning efficacy.
We then demonstrate the contribution of applying the TCP module to both the video-level branch and the moment-level branch. As depicted in Table 6, leveraging TCP solely on the video-level branch (Row 2) improves SumR to 79.7, primarily enhancing fine-grained temporal modeling. Besides, applying TCP only to the moment-level branch (Row 3) achieves an improved SumR of 80.2, as moment-level coherence learning encourages the encoder to reason coarse-grained temporal relationships. Furthermore, the full setup (Row 4) integrates these advantages, achieving the best performance across all metrics. This dual-branch design ensures holistic temporal coherence modeling at multiple granularities, which is essential for robust temporal structure learning.
4.3.4. Base Model
To demonstrate the generalization of our method, we further integrate it into two existing PRVR models: MS-SL (Dong et al., 2022a) and GMMFormer (Wang et al., 2024b). Concretely, we directly replace the base model with these models. As shown in Table 7, our approach acts as a plug-and-play enhancement to these models on the Charades-STA dataset. Consistent performance improvements are observed across both models. The universal gains also highlight that our method is model-agnostic and effectively constructs a more comprehensive cross-modal semantic space.
4.4. Qualitative Results
4.4.1. Retrieval Results
@@
To qualitatively validate the effectiveness of our method, we report visualization comparisons of retrieval results between our method, MS-SL (Dong et al., 2022a), and GMMFormer (Wang et al., 2024b) on ActivityNet Captions in Figure 5. In the first case, MS-SL and GMMFormer superficially match videos with the action riding and mistakenly associate riding with bikes, ignoring key semantics including pushes a rake and scooter. In contrast, our method accurately retrieves the target video with correct semantics. The remaining example also shows a similar phenomenon, which demonstrates the robust and superior retrieval capability of our method.
4.4.2. t-SNE Visualization
@@
We report t-SNE visualization (van der Maaten and Hinton, 2008) to analyze the alignment of video and text features in the semantic space. For clearer observation, we randomly sample a subset of paired videos and text queries from Charades-STA. As shown in Figure 6 (a), in the base model, the distributions of video and text features exhibit limited overlap in the shared cross-modal space, indicating weak alignment between the two modalities. In contrast, our method achieves a more discriminative feature distribution. As in Figure 6 (b), video and text features are tightly interleaved in regions and the regions maintain sufficient separation to preserve semantics-specific distinctions. This suggests that our approach successfully aligns the two modalities while retaining their unique characteristics, leading to improved semantic alignment for PRVR.
4.4.3. Text-Moment Similarity
@@
To verify the ability of our method to distinguish redundant moments and query-relevant moments, we present text-moment similarity examples on TVR. As shown in Figure 7, the base model produces similarity scores with limited fluctuations, indicating weak distinction between distinct moment events. Conversely, our full model exhibits clear peak values, indicating enhanced discriminative ability in moment features. It suggests that our method effectively improves the ability to distinguish between query-relevant and query-irrelevant moments.
5. Conclusion
@@
In this paper, we present a novel framework for Partially Relevant Video Retrieval, which systematically exploits the cross-model dual nature of video-text modalities: inter-sample correlation and intra-sample redundancy. Our approach introduces three key modules. The Inter Correlation Enhancement module strengthens cross-modal alignment by identifying and leveraging semantically similar yet unpaired text-video samples as pseudo-positive pairs. The Intra Redundancy Mining module improves discriminative learning by distinguishing redundant moments and query-relevant moments, forcing the model to focus on query-relevant visual semantics. The Temporal Coherence Prediction module enhances temporal structure learning through a self-supervised video sequence prediction task. Extensive experiments on three datasets demonstrate that our framework achieves state-of-the-art performance.
Appendix A Dataset Details
We summarize dataset statistics in Table 8. Our method achieves state-of-the-art performance across these datasets with varying video durations (both short and long) and distinct data distributions. It proves the strong generalization capability across diverse temporal scales, moment distributions, and temporal granularities.
Appendix B Implementation Details
For video and text representations, we employ the same features provided by (Dong et al., 2022a) to ensure a fair comparison with previous works on PRVR (Dong et al., 2022a; Jiang et al., 2023; Wang et al., 2024b; Yin et al., 2024; Dong et al., 2023; Zhang et al., 2025; Moon et al., 2025; Cho et al., 2025; Jun et al., 2025). In particular, on TVR, the 3,072-D visual features are provided by (Lei et al., 2020) using concatenated frame-level ResNet152 (He et al., 2016) feature and segment-level I3D (Carreira and Zisserman, 2017) feature, while the 768-D text features are provided by (Lei et al., 2020) using RoBERTa (Liu et al., 2019). On ActivityNet Captions and Charades-STA, the 102,4-D visual features are provided by (Zhang et al., 2020a) and (Mun et al., 2020) using I3D, respectively. The 1,024-D text features are provided by (Dong et al., 2022a) using RoBERTa.
For the Transformer encoders, each encoder consists of two standard Transformer encoder layers (Vaswani et al., 2017) with four attention heads. During training, dropout (Srivastava et al., 2014) is set to 0.15 to prevent overfitting.
Appendix C Training Loss Details
Our proposed method is optimized by the loss of each module:
(12) $\mathcal{L}=\mathcal{L}_{base}+\lambda_{ice}\mathcal{L}_{ice}+\lambda_{irm}\mathcal{L}_{irm}+\lambda_{tcp}\mathcal{L}_{tcp},$
where $\lambda_{ice}$, $\lambda_{irm}$, and $\lambda_{tcp}$ are hyperparameters to balance losses. During experiments, $\lambda_{ice}$, $\lambda_{irm}$, and $\lambda_{tcp}$ are set to 0.1, 1, and 1 on all datasets, respectively. We present ablations on these loss weights in Section D.4.
Table 8. Statistics of datasets used in our experiments.
Dataset Average Length (s) Moment-to-Video Ratio
moments videos min max mean
TVR 9.1 76.2 0.48% 100% 11.9%
ActivityNet Captions 36.2 117.6 0.48% 100% 30.8%
Charades-STA 8.1 30.0 4.3% 100% 26.3%
Table 9. Ablation study of the video-level branch and the moment-level branch in the base model on Charades-STA.
Row Setting R@1 R@5 R@10 R@100 SumR
Video Moment
1 2.4 7.3 13.5 51.4 74.6
2 2.6 9.0 14.5 53.3 79.4
3 2.9 9.2 14.9 54.3 81.3
Table 10. Ablation study of the contrastive learning losses on Charades-STA.
Row Setting R@1 R@5 R@10 R@100 SumR
Triplet Ranking InfoNCE
1 2.2 7.9 14.1 52.2 76.4
2 2.3 7.6 12.9 49.8 72.6
3 2.9 9.2 14.9 54.3 81.3
Table 11. Performance of long video retrieval on ActivityNet Captions.
Setting R1 R5 R10 R100 SumR
All queries 10.1 28.6 41.9 79.8 160.4
Queries paired with videos ¿ 3 mins 10.5 28.5 42.0 80.1 161.1
Table 12. Ablation study of $\lambda_{ice}$ across all three datasets on the metric SumR.
$\lambda_{ice}$ Charades-STA TVR ActivityNet Captions
0.05 80.0 191.2 157.1
0.1 81.3 194.0 160.4
0.5 79.2 188.1 154.1
1 77.7 184.2 151.1
Table 13. Ablation study of $\lambda_{irm}$ across all three datasets on the metric SumR.
$\lambda_{irm}$ Charades-STA TVR ActivityNet Captions
0.5 78.4 187.2 153.0
1 81.3 194.0 160.4
2 79.2 190.7 155.2
4 78.3 186.2 153.3
Table 14. Ablation study of $\lambda_{tcp}$ across all three datasets on the metric SumR.
$\lambda_{tcp}$ Charades-STA TVR ActivityNet Captions
0.5 80.2 192.4 161.7
1 81.3 194.0 160.4
2 78.3 188.7 154.9
4 76.2 183.2 150.1
Table 15. Ablation study of the correlation threshold in ICE across all three datasets on the metric SumR.
Threshold Charades-STA TVR ActivityNet Captions
0.2 80.1 188.4 156.7
0.4 81.3 194.0 160.4
0.6 79.6 189.7 156.9
0.8 78.6 186.1 155.8
Appendix D More Experiments
D.1. Multi-scale Branches
We validate the effectiveness of the video-level and moment-level branches in the base model. As shown in Table 9, combining both video-level and moment-level branches (Row 3) achieves superior performance compared to using either branch alone (Row 1 and 2). This combined setup outperforms both individual branches, demonstrating that global and local temporal cues are complementary.
D.2. Contrastive Learning Losses
We verify the contributions of the contrastive learning losses. As summarized in Table 10, combining triplet ranking loss and InfoNCE loss (Row 3) achieves the best performance, demonstrating their complementary roles in optimizing the cross-modal semantic space.
D.3. Long Video Retrieval
We further evaluate retrieval performance on long videos (¿3 minutes). While Charades-STA and TVR contain limited long-video samples in test sets (0 and 1, respectively), ActivityNet Captions provides 1,019 test videos (4,039 moment-query pairs) exceeding 3 minutes. Note that we maintain the same number of video candidates (i.e., all videos in the test set) for fair comparison. As shown in Table 11, the retrieval accuracy for text queries paired with these long videos remains comparable with the overall performance, verifying our method’s robustness across varying video durations.
D.4. Loss Weights
We present a comprehensive analysis of the loss weights. The ablations on $\lambda_{ice}$, $\lambda_{irm}$, and $\lambda_{tcp}$ are provided in Table 12, 13, and 14, respectively. It is worth noting that our loss weight selections achieve nearly optimal performance across all datasets, indicating strong generalization abilities under diverse scenarios.
Table 16. Ablation study on the redundant features in IRM across all three datasets on the metric SumR.
Row $r_{v}$ $r_{q}$ Charades-STA TVR ActivityNet Captions
1 76.8 184.7 149.9
2 77.9 186.1 154.7
3 78.7 187.8 156.1
4 81.3 194.0 160.4
Table 17. Ablation study of the shuffle ratio in TCP across all three datasets on the metric SumR.
Ratio Charades-STA TVR ActivityNet Captions
0.15 80.3 190.4 156.9
0.25 81.3 194.0 160.4
0.5 79.4 187.6 154.2
0.75 78.2 184.9 153.4
1.00 77.7 183.7 151.8
Table 18. Ablation study on methods in generating video moments across all three datasets on the metric SumR.
Method TVR ActivityNet Captions Charades-STA
Multi-scale Sliding Windows (Dong et al., 2022a) 194.6 160.1 81.6
Fixed-scale Sampling (Ours) 194.0 160.4 81.3
D.5. Correlation Threshold
We validate the efficacy of the correlation threshold in the Inter Correlation Enhancement (ICE) module. As presented in Table 15, the choice of the threshold affects retrieval performance. When it is set to 0.4, the model achieves the highest SumR across all datasets, outperforming both lower and higher thresholds. Lowering the threshold to 0.2 increases noise from weakly correlated pairs, reducing SumR on all datasets. On the other hand, raising it to 0.6 or 0.8 over-suppresses potential correlations, thus degrading SumR.
D.6. Redundant Features
We assess the effectiveness of each redundant feature in the Intra Redundancy Mining (IRM) module. As shown in Table 16, both $r_{v}$ and both $r_{q}$ alone lead to performance boosts in SumR on all datasets (Row 2 and 3). Moreover, their combination leads to further improvements (Row 4), indicating their complementary nature in enhancing visual discrimination.
D.7. Shuffle Ratio
We assess the impact of the shuffle ratio in the Temporal Coherence Prediction (TCP) module. As indicated in Table 17, the choice of the ratio influences the model’s ability to learn temporal coherence. When it is set to 0.25, the model achieves the optimal SumR on all datasets, outperforming all other configurations. Adopting other ratios results in degradations in performance: smaller ratios yield insufficient learning signals, as minimal shuffling fails to challenge the model’s temporal reasoning. Conversely, larger ratios excessively disrupt temporal coherence, making the reordering task overly ambiguous and harming feature learning.
D.8. Methods in Generating Video Moments
We verify the robustness of our method in the way of generating video moments. We compare with another moment generation strategy (Dong et al., 2022a), which employs multi-scale sliding windows to extract moments. As presented in Table 18, different generation methods only bring a slight result shift. Since we fully exploit the rich semantics of fine-grained moments and improve discrimination of visual semantics within videos, our model can directly extract discriminative moment-level features with diverse semantics, rather than depending on explicit predefined segment boundaries to extract all possible moments.

### v4

@@ 4.1.2 Evaluation Metrics
Following previous works [MSSL, DKD, GMMFORMER], we adopt the rank-based metrics Recall Rate at $k$ (R@$k$) and the sum of all recall rates (SumR). R@$k$ measures the proportion of relevant items successfully retrieved within the top-$k$ results, with $k$ set to {1, 5, 10, 100}. Higher recall rates indicate better retrieval accuracy.
4.1.3 Implementation Details
@@
For fair comparison with prior works on PRVR [MSSL, PEAN, GMMFORMER, yin2024exploiting, DKD, MGAKD, PBU, HLFormer], we utilize the same visual and textual features provided by [MSSL]. The visual and textual features are extracted by pre-trained I3D [carreira2017quo] with ResNet152 [he2016deep] and RoBERTa [roberta]. The feature dimension $D$ is set to 384. $\lambda_{1}$, $\lambda_{2}$, $T_{m}$, and $\alpha$ are set to 0.02, 0.04, 32, and 0.7 following [MSSL, GMMFORMER]. $g$ is set to 8. For training, we utilize the Adam optimizer [KingmaB14] with an initial learning rate of 2.5e-4 and a mini-batch size of 128. The model is trained for 100 epochs with an early stopping strategy. All experiments are conducted on a single NVIDIA RTX 3090 GPU.
4.2 Comparison with State-of-the-Arts
4.2.1 Compared Methods
We compare with other methods on TVR, ActivityNet Captions, and Charades-STA. These methods include PRVR [MSSL, PEAN, GMMFORMER, yin2024exploiting, DKD, MGAKD, PBU, HLFormer, SDM, RAL], T2VR [wu2023cap4video, li2023unmasked], and large multimodal models [wang2024internvideo2]. For fair comparison, all compared PRVR methods adopt the same I3D/ResNet/RoBERTa features provided by [MSSL].
4.2.2 Performance
@@
As summarized in Table 1, our method achieves state-of-the-art (SOTA) results on all three datasets across all metrics. Notably, we outperform the previous SOTA method MGAKD [MGAKD] without using additional CLIP models. Besides, we also surpass the Large Multimodal Model (LMM) InternVideo2 [wang2024internvideo2] by a significant margin. InternVideo2 merely considers the similarity between text queries and entire videos, neglecting the essential fine-grained understanding of moment events. In contrast, our method fully leverages the rich semantics of distinct moment events and learns to distinguish them. Comparison highlights the advantages and expertise of our method in the PRVR task.
To demonstrate the generalization of our method, we also integrate it into two existing PRVR models: MS-SL [MSSL] and GMMFormer [GMMFORMER]. Concretely, we directly replace the base model with these models. As shown in Table 1, our approach acts as a plug-and-play enhancement to these models on all datasets. The universal gains also highlight that our method is model-agnostic and effectively constructs a more comprehensive cross-modal semantic space.
4.2.3 Efficiency
@@
We evaluate the efficiency of our proposed model and existing models using official codes. For fair comparison, all experiments are conducted using the same batch size and a single RTX 3090 GPU on ActivityNet Captions. All models are equipped with early stopping. As shown in Table 2, with the same I3D/RoBERTa features and fewer learnable parameters, we achieve faster training/inference speed and better retrieval accuracy. These results demonstrate the effectiveness and efficiency of our proposed model.
@@
Table 2: Efficiency comparison on ActivityNet Captions using the same I3D and RoBERTa features under identical conditions.
Method Model Size Training Time Inference Time SumR
MS-SL [MSSL] 3.37M 166.05 min 146.63s 140.1
GMMFormer [GMMFORMER] 11.37M 89.05 min 87.69s 146.0
SDM [SDM] 25.73M 274.38 min 91.85s 156.6
HLFormer [HLFormer] 28.43M 189.12 min 85.39s 154.9
Ours 4.65M 80.56 min 83.01s 160.4
4.3 Ablation Studies
4.3.1 Main Ablation Studies
Table 3: Main ablation studies of the proposed key modules.
Row Setting TVR ActivityNet Captions Charades-STA
ICE IRM TCP R@1 R@5 R@1 R@5 R@1 R@5
1 13.8 32.2 7.0 22.9 1.5 6.8
2 14.6 33.9 7.8 24.9 2.2 7.5
3 14.4 33.7 7.7 24.5 2.1 7.8
4 14.1 33.5 7.4 24.1 1.7 7.0
5 15.7 36.5 8.1 26.8 2.4 8.3
6 15.4 36.2 8.3 26.4 2.5 8.5
7 16.0 36.9 9.1 27.0 2.6 8.8
8 17.5 39.0 10.1 28.6 2.9 9.2
Table 4: Ablation study of loss terms in the IRM module.
Row Loss Terms TVR ActivityNet Captions Charades-STA
$\mathcal{L}_{neg}$ $\mathcal{L}_{red}$ R@1 R@5 R@1 R@5 R@1 R@5
1 15.4 36.2 8.3 26.4 2.5 8.5
2 16.4 38.1 9.2 27.5 2.7 8.8
3 15.9 37.1 9.1 27.8 2.6 8.8
4 17.5 39.0 10.1 28.6 2.9 9.2
@@ 4.3.1 Main Ablation Studies
(1) Each module independently improves performance over the base model (Row 1-4), suggesting their effectiveness in refining the cross-modal semantic space.
@@
(3) Integrating all three modules leads to the best performance (Row 8), with significant gains across all metrics. In particular, TCP further improves ICE and IRM. ICE and IRM improve R@1 by 2.2 on TVR (Row 1 and 7, 13.8 → 16.0). When combined with TCP, the performance gain increases to 3.4 (Row 4 and 8, 14.1 → 17.5).
4.3.2 Intra Redundancy Mining
@@
We verify the effectiveness of each loss function in the Intra Redundancy Mining (IRM) module. As shown in Table 4, both $\mathcal{L}_{neg}$ (Row 2) and $\mathcal{L}_{red}$ (Row 3) lead to boosts on all metrics. Moreover, combining $\mathcal{L}_{neg}$ and $\mathcal{L}_{red}$ (Row 4) yields the highest results. The effectiveness arises because $\mathcal{L}_{neg}$ directly suppresses redundancy by pushing away redundant and query-relevant moments, while $\mathcal{L}_{red}$ implicitly strengthens cross-modal alignment by correlating video-view and query-view redundancies.
Table 5: Ablations of the group number $g$ in the TCP module.
Row $g$ TVR ActivityNet Captions Charades-STA
R@1 R@5 R@1 R@5 R@1 R@5
1 4 17.1 38.4 9.7 28.1 2.7 9.0
2 8 17.5 39.0 10.1 28.6 2.9 9.2
3 16 16.4 37.2 9.2 27.4 2.4 8.6
4 32 15.7 36.8 8.4 26.7 2.2 7.9
@@
Figure 5: Visualization comparisons of retrieval results between our method, MS-SL [MSSL], and GMMFormer [GMMFORMER] on ActivityNet Captions. The top-5 retrieval results are shown from left to right. Ground-truth videos are marked with red boxes. Zoom in for better visibility.
((a))
((b))
@@
Figure 6: t-SNE visualization [tsne] of text and video features on Charades-STA. (a) is the base model trained with only $\mathcal{L}_{base}$. (b) shows the full model trained with the full setup.
((a))
((b))
((c))
@@
Figure 7: Text-moment similarity on TVR. We smooth out similarity intervals for better readability.
4.3.3 Temporal Coherence Prediction
@@
We assess the group number $g$ in the Temporal Coherence Prediction (TCP) module, which controls the granularity of temporal coherence learning. As shown in Table 5, performances decline as $g$ >> 8 in all datasets due to increased task difficulty. The optimal results are achieved at $g$ = 8, suggesting that moderate grouping balances task difficulty and learning efficacy.
4.4 Qualitative Results
4.4.1 Retrieval Results
@@
To qualitatively validate the effectiveness of our method, we report visualization comparisons of retrieval results between our method, MS-SL [MSSL], and GMMFormer [GMMFORMER] on ActivityNet Captions in Figure 5. In the first case, MS-SL and GMMFormer superficially match videos with the action riding and mistakenly associate riding with bikes, ignoring key semantics including pushes a rake and scooter. In contrast, our method accurately retrieves the target video with correct semantics. The remaining example also shows a similar phenomenon, which demonstrates the robust and superior retrieval capability of our method.
4.4.2 t-SNE Visualization
@@
We report t-SNE visualization [tsne] to analyze the alignment of video and text features in the semantic space. For clearer observation, we randomly sample a subset of paired videos and text queries from Charades-STA. As shown in Figure 6 (a), in the base model, the distributions of video and text features exhibit limited overlap in the shared cross-modal space, indicating weak alignment between the two modalities. In contrast, our method achieves a more discriminative feature distribution. As illustrated in Figure 6 (b), video and text features are tightly interleaved in regions, and the regions maintain sufficient separation to preserve semantics-specific distinctions. This suggests that our approach successfully aligns the two modalities while retaining their unique characteristics, leading to improved semantic alignment for PRVR.
4.4.3 Text-Moment Similarity
@@
To verify the ability of our method to distinguish redundant moments and query-relevant moments, we present text-moment similarity examples on TVR. As shown in Figure 7, the base model produces similarity scores with limited fluctuations, indicating the weak distinction between distinct moment events. Conversely, our full model exhibits clear peak values, indicating enhanced discriminative ability in moment features. It suggests that our method effectively improves the ability to distinguish between query-relevant and query-irrelevant moments.
5 Conclusion
@@
In this paper, we present a novel framework for Partially Relevant Video Retrieval, which systematically exploits the cross-model dual nature of video-text modalities: inter-sample correlation and intra-sample redundancy. Our approach introduces three key modules. The Inter Correlation Enhancement module strengthens cross-modal alignment by identifying and leveraging semantically similar yet unpaired text-video samples as pseudo-positive pairs. The Intra Redundancy Mining module improves discriminative learning by distinguishing redundant moments and query-relevant moments, forcing the model to focus on query-relevant visual semantics. The Temporal Coherence Prediction module enhances discrimination of fine-grained moment-level semantics through a self-supervised video sequence prediction task. Extensive experiments on three datasets demonstrate that our framework achieves state-of-the-art performance.